OpenStudy (anonymous):

Which is an example of the associative property? A. If x = a, and a = 7, then x = 7. B. x + 5 = 5 + x C. (x + 3) + 7 = x + (3 + 7) D. 3 • (x + 5) = 3 • x + 15 I don't want the answer I am hoping you can walk me through it... Thanks! :)

OpenStudy (anonymous):

Associative property is (3 x 2) x 4 = 24 or 3 x (2 x 4) = 24. (2 + 5) + 4 = 11 or 2 + (5 + 4) = 11 Basically, you switch the parentheses around but it doesn't change the answer. :)

OpenStudy (anonymous):

I am really lost on associative property

OpenStudy (anonymous):

The answer is C. http://www.mathsisfun.com/associative-commutative-distributive.html

OpenStudy (anonymous):

You would put the parentheses on the other side of the equation

OpenStudy (anonymous):

@MiaKitty1 is right.

OpenStudy (anonymous):

Like this one.. (3 x 2) x 4 = 24 becomes 3 x (2 x 4) = 24.

OpenStudy (anonymous):

So switch up the equation and it equals the same?

OpenStudy (anonymous):

Yes, that's right

OpenStudy (anonymous):

How will I know how to switch it? Is it something you just keep triyng

OpenStudy (anonymous):

I mean it normally would ask you to do associative property if they need you to do this. :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

to me associative just means order doesn't matter like a+b+c = a+ c+b